# Learning Unknown Intervention Targets in Structural Causal Models from Heterogeneous Data

### 11 Dec 2023 v1

@@ D.4 Proof of Theorem 1
D.5 Proof of Theorem 2
To better distinguish between intervention targets in ${\bf T_{O}}$ and in ${\bf T_{L}}$, we use $\tilde{N}_{h_{i}}$ to represent the exogenous noise corresponding to $X_{i}$ for all $X_{i}\in{\bf T_{L}}$. For ease of notation, we denote the true noises $N_{j}$ in the augmented graph as $\tilde{N}_{j}$ in the following proof, since there is a one-to-one correspondence between them. We first present the following remark:
@@ D.5 Proof of Theorem 2
If a group of variables $\mathcal{X}_{i}=\{X_{i_{1}},\cdots,X_{i_{p}}\}$ have the same indicator set and $\mathcal{N}_{i}\neq\emptyset$ (note that $\mathcal{N}_{i_{k}}$ is the same for all $k\in[p]$), then any root variable in $\mathcal{X}_{i}$ (i.e., variable with no parent in $\mathcal{X}_{i}$) must be a child of all recovered noises in $\mathcal{N}_{i_{1}}$.
D.5.1 Proof of sufficiency
@@ D.6 Proof of Remark 2
Please note that $\mathcal{N}_{i}$ only contains $\tilde{N}_{l}$ under causal sufficiency assumption (otherwise, at least two variables are descendants of each other which is impossible) and each recovered noise only has one child (i.e., their corresponding intervention target) and cannot block any path. Moreover, for any $X_{i_{j}}$, the set $K=\{i_{k}\}$ and ${\bf S}={\bf S}_{i_{1}}$ is enough to guarantee that $\tilde{N}_{l}\perp\!\!\!\perp X_{i_{j}}|X_{i_{k}}\cup{\bf S}_{i_{1}}$ where $X_{i_{k}}\in{\bf T}$ and $\tilde{N}_{l}$ is its corresponding recovered noise. Thus, all $X_{i_{j}}$ with $j\neq k$ are excluded from ${\bf K}$. Moreover, $X_{i_{k}}$ is added to ${\bf K}$ as $\tilde{N}_{l}$ is the direct parent of $X_{i_{k}}$ and cannot be d-separated by a subset of observed variables and recovered noises.
@@
About comparing our result with (Jaber et al.,, 2020) In the case that ${\bf T}_{{\bf O}}={\bf T}$, the augmented graph in (Jaber et al.,, 2020) is constructed as follows (which we show it by $G^{\prime}_{{\bf T}}$). For any pairs of environments such as $E_{i}$ and $E_{j}$, a node denoted by $F_{ij}$ is added to the original graph, and it is connected to all variables in ${\bf T}_{{\bf O}}$. Please note that in our setting, all variables in ${\bf T}$ are changing across the environments and all $F_{ij}$s are directly connected to ${\bf T}_{{\bf O}}$. The candidate set in (Jaber et al.,, 2020) is the set of neighbors of $F$-nodes in the MAG of the augmented graph. Now, based on Theorem 2, it is just needed to show that if some $\tilde{N}_{i}$ is a parent of some $X_{j}$ in $Aux(G_{\bf T})$, then some $F$-node is also connected to $X_{j}$ in MAG of $G^{\prime}_{{\bf T}}$. If $X_{j}$ is ${\bf T}_{{\bf O}}$, this is trivial as it should be connected to some recovered noise in $Aux(G_{\bf T})$ and also to some $F$-node in the MAG of $G^{\prime}_{{\bf T}}$. Thus, suppose that the $X_{i}\not\in{\bf T}_{{\bf O}}$. In that case, based on part (a)(i) of Definition 1, there should be an inducing path between $\tilde{N}_{i}$ and $X_{j}$ relative to ${\bf L}$ in $G_{{\bf T}}$. This path starts from $\tilde{N}_{i}$ and goes directly to its corresponding observed variable in ${\bf T}_{{\bf O}}$ (without loss of generality, let us say $X_{i}$) and continues until getting to $X_{j}$ in the augmented graph $G_{\bf T}$. Note that since there are no changing latent variables, this path only involves variables in ${\bf O}$ and ${\bf L}$. Therefore we can construct the same path on the original graph $G$. We denote the part of this path in the original graph by $P$. Now, in $G^{\prime}_{{\bf T}}$, due to symmetry among $F$-nodes, consider any of them. This node is connected to $X_{i}$ as $X_{i}\in{\bf T}$. Now, the path starting from the $F$-node and then going directly to $X_{i}$ and continuing based on $P$ until reaching $X_{j}$ is also an inducing path in $G^{\prime}_{{\bf T}}$. This shows that our identifiability result is stronger than the one in (Jaber et al.,, 2020) when ${\bf T}_{{\bf O}}={\bf T}$, because of part (a)(ii) in Definition 1.
@@ E.1 Simulation settings
Data Generating Mechanism. We evaluated the performance of LIT algorithm on randomly generated synthetic models. Specifically, we considered each of the three settings below for data generating mechanism:
@@
We note that FCI-JCI123 algorithm is executable only under the first two settings with $D=8$ due to huge runtime. Therefore we only report FCI-JCI123 algorithm in these two setups. We fine-tuned all three algorithms with different confidence thresholds (parameters), and report the performance of both algorithms with the best performance (PreDITEr with threshold 0.05, UT-IGSP with threshold 0.005, and FCI-JCI123 with threshold 0.2).
@@
Additionally, to evaluate how many CI/invariance tests can be reduced by LIT algorithm, we compared the number of CI tests conducted by LIT algorithm, and the number of precision difference estimations (PDE estimates) in PreDITEr algorithm, which are used to replace invariance tests in their algorithm (Varici et al.,, 2022). We note that in linear SCMs (i.e., settings 1 and 3), as explained in Appendix C, the indicator set $\mathcal{I}$ can be deduced from the recovered mixing matrix from linear ICA. Therefore we only reported the number of CI tests in (C1) (resp. (C2)) in condition 5 (resp. C2) for both settings.
@@ E.2 Simulation results
Linear Gaussian model under causal sufficiency assumption. We selected $n=[5,6,7,8,9,10,$ 11,12]11,12]11 , 12 ] and $D=[8,16,32]$. We repeated the simulation for 40 times for each $n$ and $D$. We calculated the mean and the errorbar of the metrics, where the errorbar is defined as the 25% and 75% percentile among all 40 samples. The plots are shown in Figure 4. The rows represent different number of environments $D$ (32, 24, 8), and the columns represent different metrics (F1 score, precision, recall). We observe that PreDITEr algorithm has the best performance, where the $F1$ score is almost one. We note that this is because PreDITEr algorithm is designed specifically for intervention target estimation in linear Gaussian SCMs, which utilizes the sample covariance matrix in the recovery. Besides, UT-IGSP algorithm and LIT algorithm also have decent performances in this setup. Lastly, we observe that the performance of FCI-JCI123 is significantly worse than the other three methods. We note that the simulation results in Varici et al., (2022); Mooij et al., (2020) show that FCI-JCI has comparable performance when the intervention affects the noise distribution to have a mean shift. However, in our setting, the intervention only affects the noise distribution to have a changing variance (scale).
Nonlinear model under causal sufficiency assumption. We selected $n=[5,6,7,8,9,10,$ 11,12]11,12]11 , 12 ] and $D=[8,16,32]$, and also repeated the simulation for 40 times for each $n$ and $D$. The average of the metrics are shown in Figure 5. We observe that the performance of PreDITEr algorithm is worse than in Figure 4. This implies that the performance of PreDITEr algorithm depends heavily on the linear Gaussian assumption of the underlying model. Further, LIT algorithm and UT-IGSP algorithms have similar performance, where LIT performs better when the number of environment is large. This is because the number of environments has a significant impact on the recovery of the noises in nonlinear ICA. Lastly, FCI-JCI123 still have the worst performance.
@@
We observe that LIT algorithm has the best performance among all three algorithms. In particular, all three algorithms have high value in recall, and LIT algorithm has the highest precision. This means that all three algorithms can recover the intervention target in ${\bf T_{O}}$. However, since UT-IGSP algorithm does not allow for any latent confounders, and PreDITEr algorithm only allows for latent confounders that are not in ${\bf T}$, both algorithms cannot distinguish between a changing latent confounder and a changing observed variable. On the contrary, LIT algorithm can detect the changing latent confounders and thus have a better recovery of the ${\bf T_{O}}$. Lastly, note that in this setting, the average F1-score of both algorithm is much less than one. This is because there are intervention targets that cannot be uniquely recovered, such as the example described in Example 1.
Number of conducted CI tests. We recorded the average number of CI tests conducted by our algorithm, and the average number of PDE estimates by PreDITEr algorithm. The results for $D=32$ and $D=16$ are shown in Table 1 and 2, respectively. The results show that LIT algorithm significantly reduces the number of CI tests conducted, especially when $D$ is large. Specifically, the total number of CI tests by LIT algorithm grows quadratically as $n$ increases and is less than 80, while the total number of PDE estimates depends can reach up to 28,900.
Figure 4: Performance of the algorithms under Setting (1) (Linear Gaussian model under causal sufficiency assumption). Error bars represents the 25% and 75% percentiles of the corresponding metrics. Note that for $D=32$, PreDITEr algorithm (orange line) overlaps with UT-IGSP algorithm (green line) in recall, as both algorithms have 1.0 recall for all $n$.
@@
Figure 6: Performance of the algorithms under Setting (3) (Linear Gaussian model in the presence of latent confounders). Error bars represents the 25% and 75% percentiles of the corresponding metrics. Note that PreDITEr algorithm (orange line) overlaps with UT-IGSP algorithm (green line) in recall for $D=32$ and $D=16$. Further, the are no error bars for recall, because the 25% percentile is the same as the 75% percentile (i.e., 1.0) for all three algorithms.
Table 1: Average number of CI tests in LIT algorithm / PDE estimates in PreDITEr algorithm for $D=32$.
@@
LIT 1.05 0.80 1.15 1.00 1.30 1.15 1.00
PreDITEr 1460.73 1941.05 2239.43 2275.75 2764.68 3327.98 4076.13
Table 2: Average number of CI tests in LIT algorithm / PDE estimates in PreDITEr algorithm for $D=16$.

### 09 Mar 2024 v2

@@ D.4 Proof of Theorem 1
D.5 Proof of Proposition 3
The proof of the three statements in the proposition are as follows.
I The same proof of (I) in Proposition 2 in Appendix D.3 can be applied here.
IV By contradiction, suppose $X_{i}\in{\bf T_{O}}$. Then its corresponding exogenous noise $\tilde{N}_{i}$ must be in $\mathcal{N}_{i}$. There are two cases: (1) $\tilde{N}_{i}$ does not belong to any other indicator set. (2) $\tilde{N}_{i}$ belongs to another indicator set ${\bf I}_{j_{l}}$ for some $X_{j_{l}}\neq X_{i}$. Then $X_{j_{l}}$ must be a descendant of $X_{i}$, hence $\mathbf{I}_{i}\subseteq\mathbf{I}_{j_{l}}$ under Assumption 2. Therefore both cases contradict the condition in (IV).
C2 Similar to the proof of 5 in Appendix D.3, each noise in $\mathcal{N}_{i}$ corresponds to either a latent confounder, or an observed variable in $\mathcal{X}_{i}$. Further, at most one noise corresponds to a variable in $\mathcal{X}_{i}$.
We first present the following remark:
@@ D.5 Proof of Proposition 3
Any root variable in $\mathcal{X}_{i}$ (i.e., variable with no parent in $\mathcal{X}_{i}$) must be a child of all recovered noises in $\mathcal{N}_{i}$ in the augmented graph $G_{\bf T}$. (Recall that in the definition of $G_{\bf T}$, all latent variables in ${\bf T_{L}}$ are replaced by their corresponding exogenous noise.)
To prove Remark 9, suppose $X_{i_{k}}$ is a root variable in $\mathcal{X}_{i}$. Then for any noise $\tilde{N}_{l}\in\mathcal{N}_{i}$, and its corresponding variable $X_{l}$, there are only two possible cases: (1) $X_{l}\in\mathcal{X}_{i}$, (2) $X_{l}\in{\bf T_{L}}$, and it is not a parent of any observed variable in ${\bf S}_{i}$. For case (1), since $X_{i_{k}}$ is a root variable, $X_{l}$ must be $X_{i_{k}}$, otherwise $X_{i_{k}}$ has a parent in $\mathcal{X}_{i}$. Similarly, for case (2), $X_{l}$ must be a parent of $X_{i_{k}}$, hence $N_{l}$ is a direct parent of $X_{j_{k}}$.
In the following we prove C2 by contradiction. Suppose $X_{i_{j}}\in{\bf T_{O}}$. Then all other variables in $\mathcal{X}_{i}$ are descendants of $X_{i_{j}}$, and $X_{i_{j}}$ is the only root variable in $\mathcal{X}_{i}$. Then according to Remark 9, for each recovered noise $\tilde{N}_{l}\in\mathcal{N}_{i}$, there will be an edge from the corresponding noise $N_{l}$ to $X_{i_{j}}$ in $G_{\bf T}$, no matter whether $X_{l}$ is an observed variable (i.e., $X_{l}=X_{i_{j}}$) or a latent confounder. This edge cannot be blocked by any other variables or noises, and it implies that Condition (C2) cannot be satisfied for $X_{i_{j}}$, which leads to contradiction.
Lastly, we describe why II does not necessarily hold any more. This is because $X_{k}$ in the proof of II may be a latent variable (i.e., in ${\bf T_{L}}$). In this case, $\mathbf{I}_{k}$ is not observable, and $X_{k}\not\in{\bf S}_{i}$. However, if we are given the prior knowledge that ${\bf T_{L}}=\emptyset$, then II still holds.
D.6 Proof of Theorem 2
To better distinguish between intervention targets in ${\bf T_{O}}$ and in ${\bf T_{L}}$, we use $\tilde{N}_{h_{i}}$ to represent the exogenous noise corresponding to $X_{i}$ for all $X_{i}\in{\bf T_{L}}$. For ease of notation, we denote the true noises $N_{j}$ in the auxilary graph as $\tilde{N}_{j}$ in the following proof, since there is a one-on-one correspondence between them.
D.6.1 Proof of Sufficiency
@@ D.7 Proof of Remark 3
Please note that $\mathcal{N}_{i}$ only contains $\tilde{N}_{l}$ (otherwise, at least two variables are descendants of each other which is impossible) under causal sufficiency assumption and each recovered noise only has one children (i.e., their corresponding intervention target) and cannot block any path. Moreover, for any $X_{i_{j}}$, the set $K=\{i_{k}\}$ and ${\bf S}={\bf S}_{i_{1}}$ is enough to guarantee that $\tilde{N}_{l}\perp\!\!\!\perp X_{i_{j}}|X_{i_{k}}\cup{\bf S}_{i_{1}}$ where $X_{i_{k}}\in{\bf T}$ and $\tilde{N}_{l}$ is its corresponding recovered noise. Thus, all $X_{i_{j}}$s, $j\neq k$ are excluded from ${\bf K}$. Moreover, $X_{i_{k}}$ is added to ${\bf K}$ as $\tilde{N}_{l}$ is the direct parent of $X_{i_{k}}$ and cannot be d-separated by a subset of observed variables and recovered noises.
@@
About comparing our result with Jaber et al., (2020) In the case that ${\bf T}_{{\bf O}}={\bf T}$, the augmented graph in Jaber et al., (2020) is constructed as follows (which we show it by $G^{\prime}_{{\bf T}}$). For any pairs of environments such as $E_{i}$ and $E_{j}$, a node denoted by $F_{ij}$ is added to the original graph, and it is connected to all variables in ${\bf T}_{{\bf O}}$. Please note that in our setting, all variables in ${\bf T}$ are changing across the environments and all $F_{ij}$s are directly connected to ${\bf T}_{{\bf O}}$. The candidate set in Jaber et al., (2020) is the neighbors of $F$-nodes in the MAG of the augmented graph. Now, based on Theorem 2, it is just needed to show that if some $\tilde{N}_{i}$ is a parent of some $X_{j}$ in $Aux(G_{\bf T})$, then some $F$-node is also connected to $X_{j}$ in MAG of $G^{\prime}_{{\bf T}}$. If $X_{j}$ is ${\bf T}_{{\bf O}}$, this is trivial as it should be connected to some recovered noise in $Aux(G_{\bf T})$ and also to some $F$-node in the MAG of $G^{\prime}_{{\bf T}}$. Thus, suppose that the $X_{i}\not\in{\bf T}_{{\bf O}}$. In that case, based on part (a)(i) of Definition 1, there should be an inducing path between $\tilde{N}_{i}$ and $X_{j}$ relative to ${\bf L}$ in $G_{{\bf T}}$. This path starts from $\tilde{N}_{i}$ and goes directly to its corresponding observed variable in ${\bf T}_{{\bf O}}$ (without loss of generality, let us say $X_{i}$) and continues until getting to $X_{j}$ in the augmented graph $G_{\bf T}$. Note that since there are no changing latent variables, this path only involves variables in ${\bf O}$ and ${\bf L}$. Therefore we can construct the same path on the original graph $G$. We denote the part of this path in the original graph by $P$. Now, in $G^{\prime}_{{\bf T}}$, due to symmetry among $F$-nodes, consider any of them. This node is connected to $X_{i}$ as $X_{i}\in{\bf T}$. Now, the path starting from the $F$-node and then going directly to $X_{i}$ and continues based on $P$ until reaching $X_{j}$ is also an inducing path in $G^{\prime}_{{\bf T}}$. This shows that our identifiability result is stronger than the one in Jaber et al., (2020) when ${\bf T}_{{\bf O}}={\bf T}$, because of part (a)(ii) in Definition 1.
@@ E.1 Simulation Settings
Data Generating Mechanism. We evaluated the performance of LIT algorithm on randomly generated synthetic models. [^7] All numerical simulations are run on the CPU of an Apple Macbook Pro laptop with M1 Pro chip. Specifically, we considered each of the three settings below for data generating mechanism:
@@
    CITE algorithm [^10] https://github.com/bvarici/intervention-estimation in (Varici et al.,, 2021). This is the prior work of PreDITEr algorithm, which only works for linear Gaussian SCMs under causal sufficiency.
  4. 4.
@@
We note that FCI-JCI123 algorithm is executable only under the first two settings with $D=8$ due to huge runtime. Therefore we only report FCI-JCI123 algorithm in these two setups. We fine-tuned all three algorithms with different confidence thresholds (parameters), and report the performance of both algorithms with the best performance (PreDITEr with threshold 0.05, UT-IGSP with threshold 0.005, CITE with threshold 0.5, and FCI-JCI123 with threshold 0.2).
@@
Additionally, to evaluate how many CI/invariance tests can be reduced by LIT algorithm, we compared the number of CI tests conducted by LIT algorithm, and the number of precision difference estimations (PDE estimates) in PreDITEr algorithm and CITE algorithm, which are used to replace invariance tests in the algorithm (Varici et al.,, 2022). We note that in linear SCMs (i.e., settings 1 and 3), as explained in Appendix C, the indicator set $\mathcal{I}$ can be deduced from the recovered mixing matrix from linear ICA. Therefore we only reported the number of CI tests in (C1) (resp. (C2)) in condition 5 (resp. C2) for both settings.
@@ E.2 Simulation Results
Linear Gaussian model under causal sufficiency assumption. We selected $n=[5,6,7,8,9,10,$ 11,12]11,12]11 , 12 ] and $D=[8,16,32]$. We repeated the simulation for 40 times for each $n$ and $D$. We calculated the mean and the errorbar of the metrics, where the errorbar is defined as the 25% and 75% percentile among all 40 samples. The plots are shown in Figure 5. The rows represent different number of environments $D$ (32, 24, 8), and the columns represent different metrics (F1 score, precision, recall). We observe that PreDITEr and CITE algorithm have the best performance, where the F1 score is almost one. We note that this is because PreDITEr and CITE algorithm are designed specifically for intervention target estimation in linear Gaussian SCMs, which utilizes the sample covariance matrix in the recovery. Besides, UT-IGSP algorithm and LIT algorithm also have decent performances in this setup. Lastly, we observe that the performance of FCI-JCI123 is significantly worse than the other three methods. We note that the simulation results in Varici et al., (2022); Mooij et al., (2020) show that FCI-JCI has comparable performance when the intervention affects the noise distribution to have a mean shift. However, in our setting, the intervention only affects the noise distribution to have a changing variance (scale).
Nonlinear model under causal sufficiency assumption. We selected $n=[5,6,7,8,9,10,$ 11,12]11,12]11 , 12 ] and $D=[8,16,32]$, and also repeated the simulation for 40 times for each $n$ and $D$. The average of the metrics are shown in Figure 6. We observe that the performance of PreDITEr algorithm is worse than in Figure 5. This implies that the performance of PreDITEr algorithm depends heavily on the linear Gaussian assumption of the underlying model. Further, LIT, CITE and UT-IGSP algorithms have similar performance, where LIT performs better when the number of environment is large. This is because the number of environments has a significant impact on the recovery of the noises in nonlinear ICA. Lastly, FCI-JCI123 still have the worst performance.
@@
We observe that LIT algorithm has the best performance among all three algorithms. In particular, all three algorithms have high value in recall, and LIT algorithm has the highest precision. This means that all three algorithms can recover the intervention target in ${\bf T_{O}}$. However, since CITE and UT-IGSP algorithm does not allow for any latent confounders, and PreDITEr algorithm only allows for latent confounders that are not in ${\bf T}$, all three algorithms cannot distinguish between a changing latent confounder and a changing observed variable. On the contrary, LIT algorithm can detect the changing latent confounders and thus have a better recovery of the ${\bf T_{O}}$. Lastly, note that in this setting, the average F1-score of both algorithm is much less than one. This is because there are intervention targets that cannot be uniquely recovered, such as the example described in Example 1.
Number of conducted CI tests. We recorded the average number of CI tests conducted by our algorithm, and the average number of PDE estimates by PreDITEr and CITE algorithm. The results for $D=32$ and $D=16$ are shown in Table 1 and 2, respectively. The results show that LIT algorithm significantly reduces the number of CI tests conducted, especially when $D$ is large. Specifically, the total number of CI tests by LIT algorithm grows quadratically as $n$ increases and is less than 80, while the total number of PDE estimates depends can reach up to 29,600.
Figure 5: Performance of the algorithms under Setting (1) (Linear Gaussian model under causal sufficiency assumption). Error bars represents the 25% and 75% percentiles of the corresponding metrics. Note that for $D=32$, PreDITEr algorithm (orange line) overlaps with UT-IGSP algorithm (green line) and CITE algorithm (red line) in recall, as both algorithms have 1.0 recall for all $n$.
@@
Figure 7: Performance of the algorithms under Setting (3) (Linear Gaussian model in the presence of latent confounders). Error bars represents the 25% and 75% percentiles of the corresponding metrics. Note that PreDITEr algorithm (orange line) overlaps with UT-IGSP algorithm (green line) and CITE algorithm (red line) in recall for $D=32$. Further, the are no error bars for recall, because the 25% percentile is the same as the 75% percentile (i.e., 1.0) for all three algorithms.
Table 1: Average number of CI tests in LIT / PDE estimates in PreDITEr and CITE for $D=32$.
@@
LIT 1.55 0.95 1.75 1.05 1.05 1.00 0.85
PreDITEr 1583.78 1636.88 2515.13 2541.40 3269.72 3291.18 3016.45
CITE 547.77 648.23 939.23 1095.97 1665.63 1823.78 2809.60
Table 2: Average number of CI tests in LIT / PDE estimates in PreDITEr and CITE for $D=16$